In analyzing the crude fiber content of a particular plant sample, 0.313N NaOH solution is required.

A. What is the mass of NaOH pellets needed to prepare 500ml of the solution? 

B. What is the M of the solution? (MW NaOH: 40g/mol)


1
Expert's answer
2021-03-17T12:21:06-0400

A. Sodium hydroxide 0.313 mol/l (0.313 N) 

Then


"0.313\ mol/l\cdot(0.5\ l)=0.1565\ mol"

"22.990+15.999+1.008=39.997"


Molar Mass of NaOH is "39.997\ g/mol".

The mass of NaOH pellets needed to prepare 500ml of the solution is


"0.1565\ mol \cdot (39.997\ g/mol)=6.260 \ g"

B.

"Molarity =M=\dfrac{\text{moles of solute}}{\text{liters of soln}}"


"M=\dfrac{0.1565\ mol}{0.5\ l}=0.313\ mol/l"